# VoiceBridge: Designing Latent Bridge Models for General Speech Restoration at Scale

### v1

@@ 5 Conclusion
Our work presents VoiceBridge, modeling diverse LQ-to-HQ tasks in GSR with a single latent-to-latent generative framework backed by a transformer architecture. By introducing three novel techniques, including scale-equivariant regularization, joint neural prior, and perceptual-aware generation process, we shape a structured latent space for LBM, uniformly reduce the distance between each LQ prior and HQ target, and align both LBM sampling and VAE decoding with human perception, allowing VoiceBridge to consistently outperform strong GSR baselines on 48 kHz benchmarks and demonstrate strong performance on unseen tasks, e.g., refining recent zero-shot speech and podcast generation methods. Ablations confirm the holistic improvements made by our innovations.
@@ A.1 Bridge-based Speech Enhancement
Schrödinger Bridge (SB) models (Schrödinger, 1932; Chen et al., 2021; Wang et al., 2021; Bunne et al., 2023) explore learning an optimal stochastic trajectory between two boundary distributions with iterative fitting, allowing for efficient and high-quality generation. Built upon SB models, tractable bridge models have been recently developed (Bunne et al., 2023), which achieves data-to-data generation process effectively exploiting the instructive information contained in the observed prior distribution. These bridge models have shown advantages over diffusion models in tasks where indicative prior information has already been provided, such as image-to-image translation (Liu et al., 2023), text-to-speech synthesis (Chen et al., 2023), image-to-video generation (Wang et al., 2024c), and speech restoration (SE) (Jukić et al., 2024).
@@ A.2 General Speech Restoration
GSR, shown by (Liu et al., 2022), focuses on the task of retrieving clean speech out of lossy recordings, such as noise addition, bandwidth limitation, dereverberation, or other acoustic degradations. Unlike the SE task which restores speech from only noise addition or another single kind of degradation, GSR pays extra focus on generating audio with high perceptual quality from acoustic scenarios with mixed degradations (Liu et al., 2022; Babaev et al., 2024).
@@ Appendix C Schrödinger Bridge Training and Inference
$\mathbf{z}_{t}$ $=\frac{\alpha_{t}\sigma_{t}^{2}}{\alpha_{s}\sigma_{s}^{2}}\mathbf{z}_{s}+\alpha_{t}\left(1-\frac{\sigma_{t}^{2}}{\sigma_{s}^{2}}\right)\mathbf{z}_{\theta}(\mathbf{z}_{s},s)+\alpha_{t}\sigma_{t}\sqrt{1-\frac{\sigma_{t}^{2}}{\sigma_{s}^{2}}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ (16)
@@ D.1 Model Architecture
VoiceBridge consists of a VAE for prior latent encoding and target latent decoding, and a transformer backbone for solving the bridge model trajectory. For the autoencoder, we utilize the Oobleck VAE (Evans et al., 2024b) architecture. The encoder and decoder are symmetric and contain 156M parameters each. The VAE compresses 48 kHz audio into 64-channel latent representations at 23.4 Hz, yielding a temporal downsampling factor of 2048. The discriminator architecture follows (Evans et al., 2024a), where we employ a multi-scale STFT discriminator with 5 separate models modeling STFT spectrograms of different window lengths. For the SB model, we adopt the transformer backbone from Stable Audio 2 (Evans et al., 2024a; b), removing the text-conditioning cross-attention layers. We set the transformer backbone to have 24 transformer layers with a hidden dimension of 1152, resulting in 544 M parameters in total.
@@ D.2 Model Training Details
The training pipeline of VoiceBridge consists of four stages: (1) pre-training a VAE with the Energy Preserving training objective $\mathcal{L}_{\text{ep-vae}}$ on clean speech; (2) fine-tuning the VAE encoder on distorted inputs with the joint neural prior loss $\mathcal{L}_{\text{np-enc}}$, while keeping the decoder frozen; (3) training a transformer-based bridge model to solve the SB in the latent space by $\mathcal{L}_{\text{bridge}}$, with both the encoder and decoder frozen; and (4) jointly fine-tuning the Bridge Transformer and the VAE decoder with $\mathcal{L}_{\text{pa-gen}}$, improving perceptual quality of generation process.
@@
$\mathcal{L}_{\text{pa-gen}}$ $=\mathcal{L}_{\text{bridge}}(\varphi)+\mathcal{L}_{\text{data}}(\mathcal{D}_{\theta}(\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})),\mathbf{x}_{0})+\mathcal{L}_{\text{hf}}(\mathcal{D}_{\theta}(\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})),\mathbf{x}_{0}),$ (23)
@@
$\mathcal{L}_{\text{hf}}=\lambda_{\text{pesq}}\mathcal{L}_{\text{pesq}}+\lambda_{\text{utmos}}\mathcal{L}_{\text{utmos}},$ (24)
@@
Appendix E Dataset Details
@@
We construct our HQ speech dataset by combining publicly available datasets: VCTK (Veaux & Yamagishi, 2017) is a multi-speaker English dataset with 109 speakers of diverse accents each reading about 400 sentences from newspaper excerpts and phonetic-rich passages, at a sampling rate of 48 kHz. HiFi-TTS (Bakhturina et al., 2021) contains speech from 10 English speakers reading LibriVox audiobooks and Project Gutenberg texts, specially filtered by a threshold of bandwidth and signal-to-noise ratio. HQ-TTS (Liu et al., 2022) is a collection of publicly available 44.1kHz or 48kHz clean speech data on the OpenSLR website, including multilingual speech from a wide range of speakers. AiShell-1 (Bu et al., 2017) and AiShell-4 (Fu et al., 2021) are two multi-speaker Mandarin datasets collected from noiseless indoor environments through different recording devices, providing non-English speech context for VoiceBridge to learn multilingual speech patterns. Bible-TTS (Meyer et al., 2022) is also a multilingual dataset, composed of speech for six languages spoken in Sub-Saharan Africa. Expresso (Nguyen et al., 2023) is a high-quality expressive speech dataset, which includes both expressively rendered read speech and improvised dialogues, offering speech in various tones and styles. EARS (Richter et al., 2024) contains anechoic speech recordings from over 100 English speakers with high demographic diversity, spanning the full range of human speech. Some of these datasets include speech samples recorded in different acoustic backgrounds, from which we select the clean audio in noiseless environments. For the datasets used for later evaluation, e.g., VCTK, the corresponding testset is removed from the training data. The combination of training data, after the filtering, totals in around 1138 hours. All recordings are resampled to a sampling rate of 48 kHz.
@@
For the noise addition and reverberation application processes of data construction, extra noise data and room impulse response (RIR) signals are required. For noise samples, we utilize the following datasets. DEMAND (Valentini-Botinhao, 2017) includes multi-channel recordings of acoustic noise in diverse environments. CHiME-2/3 (Vincent et al., 2013; Barker et al., 2015) have noise recordings taken from the two ChiME ASR challenges. TUT Urban Acoustic Scenes 2018 (Mesaros et al., 2018) is a dataset of multi-device recordings, originally used for urban acoustic scene classification. DESED (Turpault et al., 2019) is an audio dataset designed for recognizing sound event classes in domestic environments. These datasets have covered common noise in different acoustic scenarios, supporting our denoising-related tasks. Similar to the speech datasets, the testing portion is also removed for DEMAND and CHiME3, as they are later used for evaluation. Room Impulse Responses are drawn from the following datasets. SLR26/28 (Ko et al., 2017) contains simulated RIRs in different room conditions. GTU-RIR (Pekmezci & Genc, 2024) is a real RIR dataset, including the recordings of the impulse signal with different devices in different rooms. We also include other simulated RIR signals created by ourselves, following previous works.
@@
Appendix F Evaluation Details
In this section, we provide a detailed introduction to the benchmark datasets, baseline methods, and evaluation metrics that we used for evaluating VoiceBridge’s performance. We also list further evaluation settings and results on broken down degradation types in detail, and make comparison to other models.
Table 9: Dereverberation results on VCTK (Veaux & Yamagishi, 2017) convolved with simulated RIR signals with different RT60 time.
Dataset Method PESQ($\uparrow$) WVMOS($\uparrow$) UTMOS($\uparrow$) DNSMOS($\uparrow$) NISQA($\uparrow$) SRMR($\uparrow$) RT60=0.3s Voicefixer 2.321 4.124 3.533 3.138 4.369 9.044 Resemble-Enhance 2.333 4.376 3.628 3.101 4.381 9.208 UniverSE++ 2.599 4.102 3.296 3.010 4.024 9.744 VoiceBridge 2.393 4.286 3.864 3.185 4.488 9.767 RT60=0.6s Voicefixer 2.060 4.068 3.394 3.111 4.294 9.398 Resemble-Enhance 1.653 4.244 3.419 3.057 4.317 9.208 UniverSE++ 1.652 3.553 2.639 2.830 3.695 9.794 VoiceBridge 2.127 4.242 3.821 3.081 4.448 9.762 RT60=0.9s Voicefixer 1.915 4.024 3.342 3.108 4.292 9.641 Resemble-Enhance 1.394 4.139 3.232 3.067 4.353 9.207 UniverSE++ 1.407 3.251 2.329 2.755 3.602 9.604 VoiceBridge 1.970 4.195 3.757 3.058 4.358 9.919 RT60=1.2s Voicefixer 1.837 3.995 3.292 3.105 4.261 9.658 Resemble-Enhance 1.323 4.100 3.175 3.051 4.271 9.334 UniverSE++ 1.351 3.170 2.193 2.706 3.540 8.999 VoiceBridge 1.873 4.177 3.732 3.058 4.338 9.968
F.1 Evaluation datasets

### v3

@@ 5 Conclusion
Our work presents VoiceBridge, modeling diverse LQ-to-HQ tasks in GSR with a single latent-to-latent generative framework backed by a transformer architecture. By introducing three novel techniques, including scale-equivariant regularization, joint neural prior, and the denoiser to generator post-training process, we allow VoiceBridge to consistently outperform strong GSR baselines on 48 kHz benchmarks with single step inference and demonstrate strong performance on unseen tasks. Ablations confirm the holistic improvements made by our innovations.
6 Impact Statement
This paper presents work whose goal is to advance the field of Generative Bridge Models, and General Speech Restoration. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Bridge-based Speech Enhancement
Schrödinger Bridge (SB) models (Schrödinger, 1932; Chen et al., 2021; Wang et al., 2021; Bunne et al., 2023) explore learning an optimal stochastic trajectory between two boundary distributions with iterative fitting, allowing for efficient and high-quality generation. Built upon SB models, tractable bridge models have been recently developed (Bunne et al., 2023), which achieves data-to-data generation process effectively exploiting the instructive information contained in the observed prior distribution. These bridge models have shown advantages over diffusion models in tasks where indicative prior information has already been provided, such as image-to-image translation (Liu et al., 2023b), text-to-speech synthesis (Chen et al., 2023), image-to-video generation (Wang et al., 2024c), and speech restoration (SE) (Jukić et al., 2024).
@@ A.2 General Speech Restoration
GSR, shown by (Liu et al., 2022), focuses on the task of retrieving clean speech out of lossy recordings, such as noise addition, bandwidth limitation, dereverberation, or other acoustic degradations. Unlike the traditional SE task which restores speech from only noise addition or another single kind of degradation, GSR pays extra focus on generating audio with high perceptual quality from acoustic scenarios with mixed degradations (Liu et al., 2022; Babaev et al., 2024).
@@ Appendix C Schrödinger Bridge Training and Inference
$\mathbf{z}_{t}$ $=\frac{\alpha_{t}\sigma_{t}^{2}}{\alpha_{s}\sigma_{s}^{2}}\mathbf{z}_{s}+\alpha_{t}\left(1-\frac{\sigma_{t}^{2}}{\sigma_{s}^{2}}\right)\mathbf{z}_{\theta}(\mathbf{z}_{s},s)+\alpha_{t}\sigma_{t}\sqrt{1-\frac{\sigma_{t}^{2}}{\sigma_{s}^{2}}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ (18)
@@ D.1 Model Architecture
VoiceBridge consists of a VAE for prior latent encoding and target latent decoding, and a transformer backbone for solving the bridge model trajectory. For the autoencoder, we utilize the Oobleck VAE (Evans et al., 2024b) architecture. The encoder and decoder are symmetric and contain 156M parameters each. The VAE compresses 48 kHz audio into 64-channel latent representations at 23.4 Hz, yielding a temporal downsampling factor of 2048. The discriminator architecture follows (Evans et al., 2024a), where we employ a multi-scale STFT discriminator with 5 separate models modeling STFT spectrograms of different window lengths. For the SB model, we adopt the transformer backbone from Stable Audio 2 (Evans et al., 2024a, b), removing the text-conditioning cross-attention layers. We set the transformer backbone to have 24 transformer layers with a hidden dimension of 1152, resulting in 544 M parameters in total.
@@ D.2 Model Training Details
The training pipeline of VoiceBridge consists of four stages: (1) pre-training a VAE with the Energy Preserving training objective $\mathcal{L}_{\text{ep-vae}}$ on clean speech; (2) fine-tuning the VAE encoder on distorted inputs with the joint neural prior loss $\mathcal{L}_{\text{np-enc}}$, while keeping the decoder frozen; (3) training a transformer-based bridge model to solve the SB in the latent space by $\mathcal{L}_{\text{bridge}}$, with both the encoder and decoder frozen; and (4) jointly fine-tuning the Bridge Transformer and the VAE decoder with $\mathcal{L}_{\text{pt}}$, improving perceptual quality of generation process.
@@
$\mathcal{L}_{\text{pt}}$ $=\mathcal{L}_{\text{bridge}}(\varphi)+\mathcal{L}_{\text{data}}(\mathcal{D}_{\theta}(\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})),\mathbf{x}_{0})$
$+\mathcal{L}_{\text{GAN}}(\mathcal{D}_{\theta}(\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})),\mathbf{x}_{0})+\mathcal{L}_{\text{perc}}(\mathcal{D}_{\theta}(\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})),\mathbf{x}_{0}),$ (25)
@@
$\mathcal{L}_{\text{perc}}=\lambda_{\text{pesq}}\mathcal{L}_{\text{pesq}}+\lambda_{\text{utmos}}\mathcal{L}_{\text{utmos}},$ (26)
@@
Appendix E Derivations on Denoiser to Generator Post-training Objective
In this appendix, we provide a concise derivation of the optimization target of our denoiser-to-generator post-training, and explain why it changes the learning objective from conditional expectation prediction (typical of MSE-trained bridge models) to conditional distribution matching.
Problem setup.
Let the conditioning signal be
$c:=(x_{t},t,x_{1}),$
where $x_{t}$ is the intermediate state used by the bridge model, $t$ is the normalized time index, and $x_{1}$ is the LQ prior (or its latent counterpart), consistent with the main text. Let $x:=x_{0}$ denote the paired HQ target waveform. The training pairs are sampled from a joint data distribution
$p_{\text{data}}(x,c)=p(c)\,p_{\text{data}}(x\mid c).$
Our post-training refines a generator that produces an HQ estimate under condition $c$. To model a conditional distribution rather than a deterministic point estimate, we assume the generator is stochastic:
$\hat{x}=G_{\varphi}(c,\epsilon),\qquad\epsilon\sim p(\epsilon),$
which induces a conditional model distribution $p_{G}(x\mid c)$.
(A) Denoiser-to-generator post-training objective.
We use the vanilla (logistic) conditional GAN with a discriminator $D_{\omega}(x,c)\in(0,1)$ and augment it with a perceptual loss. The min–max objective is
$\min_{G}\ \max_{D}\ \Big\{$ $\mathbb{E}_{(x,c)\sim p_{\text{data}}}\big[\log D(x,c)\big]+\mathbb{E}_{c\sim p(c),\,\epsilon\sim p(\epsilon)}\big[\log(1-D(G(c,\epsilon),c))\big]$
$\quad+\ \lambda_{\text{perc}}\,\mathbb{E}_{(x,c)\sim p_{\text{data}},\,\epsilon\sim p(\epsilon)}\big[\ell_{\text{perc}}(G(c,\epsilon),x)\big]\Big\},$ (27)
where $\ell_{\text{perc}}(\hat{x},x)$ is the perceptual loss used in our method (e.g., a weighted combination of PESQ and UTMOS losses, or their differentiable surrogates). For clarity, we omit other auxiliary terms in the main text (e.g., $\mathcal{L}_{\text{data}}$ or $\mathcal{L}_{\text{bridge}}$); adding them does not change the conclusion below, since they simply introduce additional regularization on top of the distribution-matching objective.
(B) What the adversarial term optimizes (JS divergence).
Fix a generator $G$ (hence $p_{G}(\cdot\mid c)$). Under standard assumptions that $p_{\text{data}}(\cdot\mid c)$ and $p_{G}(\cdot\mid c)$ admit densities with respect to a common base measure for $p(c)$-a.e. $c$, the discriminator that maximizes the GAN value function (the first two terms in Eq. E) is given by
$D^{*}(x,c)=\frac{p_{\text{data}}(x\mid c)}{p_{\text{data}}(x\mid c)+p_{G}(x\mid c)}.$ (28)
Substituting $D^{*}$ into the GAN value function yields the classical Jensen–Shannon divergence form:
$V(D^{*},G)=-\log 4+2\,\mathbb{E}_{c\sim p(c)}\Big[\mathrm{JSD}\big(p_{\text{data}}(\cdot\mid c)\ \|\ p_{G}(\cdot\mid c)\big)\Big],$ (29)
where $\mathrm{JSD}(p\|q)$ is the Jensen–Shannon divergence. Therefore, the min–max training in Eq. E is equivalent (up to the constant $\log 4$) to minimizing
$\min_{G}\ 2\,\mathbb{E}_{c}\Big[\mathrm{JSD}\big(p_{\text{data}}(\cdot\mid c)\ \|\ p_{G}(\cdot\mid c)\big)\Big]+\lambda_{\text{perc}}\,\mathbb{E}_{(x,c),\,\epsilon}\big[\ell_{\text{perc}}(G(c,\epsilon),x)\big].$ (30)
Eq. 30 makes explicit that the adversarial component performs conditional distribution matching (via JS divergence to the data conditional), while the perceptual term encourages high subjective quality.
(C) Why this differs from expectation prediction.
Many bridge models are trained as $x_{0}$-predictors with a squared error objective (in latent or waveform space), which yields conditional expectation estimation. Concretely, consider a deterministic predictor $f(c)$ trained with
$\min_{f}\ \mathbb{E}\big[\|x-f(c)\|_{2}^{2}\big].$ (31)
The population minimizer of Eq. 31 satisfies
$f^{*}(c)=\mathbb{E}[x\mid c],$ (32)
i.e., the model learns the conditional mean (“expectation modeling”). This explains why MSE-trained $x_{0}$ predictors often produce over-smoothed outputs when the conditional distribution is multi-modal.
In contrast, the adversarial term in Eq. 30 is minimized when
$p_{G}(\cdot\mid c)=p_{\text{data}}(\cdot\mid c)\qquad\text{for }p(c)\text{-a.e. }c,$ (33)
since $\mathrm{JSD}(p\|q)\geq 0$ with equality iff $p=q$. Thus, the GAN component explicitly optimizes a distributional objective rather than a point-estimation objective.
(D) Why initializing from a bridge model helps.
Our post-training starts from a bridge model that already provides a strong conditional estimate (often close to $\mathbb{E}[x\mid c]$ under MSE-style training), and then refines it using Eq. E. Intuitively, the original bridge training provides a stable initialization that preserves the learned transport structure, while adversarial fine-tuning reduces the discrepancy between the model conditional $p_{G}(\cdot\mid c)$ and the data conditional $p_{\text{data}}(\cdot\mid c)$, and the perceptual term further steers the solutions toward human-aligned quality. Importantly, to represent a non-degenerate conditional distribution in Eq. 33, the generator must include stochasticity (e.g., the noise $\epsilon$); otherwise $p_{G}(\cdot\mid c)$ collapses to a point mass and cannot match a multi-modal $p_{\text{data}}(\cdot\mid c)$.
Appendix F Dataset Details
@@
We construct our HQ speech dataset by combining publicly available datasets: VCTK (Veaux and Yamagishi, 2017) is a multi-speaker English dataset with 109 speakers of diverse accents each reading about 400 sentences from newspaper excerpts and phonetic-rich passages, at a sampling rate of 48 kHz. HiFi-TTS (Bakhturina et al., 2021) contains speech from 10 English speakers reading LibriVox audiobooks and Project Gutenberg texts, specially filtered by a threshold of bandwidth and signal-to-noise ratio. HQ-TTS (Liu et al., 2022) is a collection of publicly available 44.1kHz or 48kHz clean speech data on the OpenSLR website, including multilingual speech from a wide range of speakers. AiShell-1 (Bu et al., 2017) and AiShell-4 (Fu et al., 2021) are two multi-speaker Mandarin datasets collected from noiseless indoor environments through different recording devices, providing non-English speech context for VoiceBridge to learn multilingual speech patterns. Bible-TTS (Meyer et al., 2022) is also a multilingual dataset, composed of speech for six languages spoken in Sub-Saharan Africa. Expresso (Nguyen et al., 2023) is a high-quality expressive speech dataset, which includes both expressively rendered read speech and improvised dialogues, offering speech in various tones and styles. EARS (Richter et al., 2024) contains anechoic speech recordings from over 100 English speakers with high demographic diversity, spanning the full range of human speech. Some of these datasets include speech samples recorded in different acoustic backgrounds, from which we select the clean audio in noiseless environments. For the datasets used for later evaluation, e.g., VCTK, the corresponding testset is removed from the training data. The combination of training data, after the filtering, totals in around 1138 hours. All recordings are resampled to a sampling rate of 48 kHz.
@@
For the noise addition and reverberation application processes of data construction, extra noise data and room impulse response (RIR) signals are required. For noise samples, we utilize the following datasets. DEMAND (Valentini-Botinhao, 2017) includes multi-channel recordings of acoustic noise in diverse environments. CHiME-2/3 (Vincent et al., 2013; Barker et al., 2015) have noise recordings taken from the two ChiME ASR challenges. TUT Urban Acoustic Scenes 2018 (Mesaros et al., 2018) is a dataset of multi-device recordings, originally used for urban acoustic scene classification. DESED (Turpault et al., 2019) is an audio dataset designed for recognizing sound event classes in domestic environments. These datasets have covered common noise in different acoustic scenarios, supporting our denoising-related tasks. Similar to the speech datasets, the testing portion is also removed for DEMAND and CHiME3, as they are later used for evaluation. Room Impulse Responses are drawn from the following datasets. SLR26/28 (Ko et al., 2017) contains simulated RIRs in different room conditions. GTU-RIR (Pekmezci and Genc, 2024) is a real RIR dataset, including the recordings of the impulse signal with different devices in different rooms. We also include other simulated RIR signals created by ourselves, following previous works.
@@
Appendix G Evaluation Details
In this section, we provide a detailed introduction to the benchmark datasets, baseline methods, and evaluation metrics that we used for evaluating VoiceBridge’s performance. We also list further evaluation settings and results on broken down degradation types in detail, and make comparison to other models. Note that results in appendix are achieved by 4 inference steps.
Table 7: Dereverberation results on VCTK (Veaux and Yamagishi, 2017) convolved with simulated RIR signals with different RT60 time.
Dataset Method PESQ($\uparrow$) WVMOS($\uparrow$) UTMOS($\uparrow$) DNSMOS($\uparrow$) NISQA($\uparrow$) SRMR($\uparrow$)
RT60=0.3s Voicefixer 2.321 4.124 3.533 3.138 4.369 9.044
Resemble-Enhance 2.333 4.376 3.628 3.101 4.381 9.208
UniverSE++ 2.599 4.102 3.296 3.010 4.024 9.744
VoiceBridge 2.393 4.286 3.864 3.185 4.488 9.767
RT60=0.6s Voicefixer 2.060 4.068 3.394 3.111 4.294 9.398
Resemble-Enhance 1.653 4.244 3.419 3.057 4.317 9.208
UniverSE++ 1.652 3.553 2.639 2.830 3.695 9.794
VoiceBridge 2.127 4.242 3.821 3.081 4.448 9.762
RT60=0.9s Voicefixer 1.915 4.024 3.342 3.108 4.292 9.641
Resemble-Enhance 1.394 4.139 3.232 3.067 4.353 9.207
UniverSE++ 1.407 3.251 2.329 2.755 3.602 9.604
VoiceBridge 1.970 4.195 3.757 3.058 4.358 9.919
RT60=1.2s Voicefixer 1.837 3.995 3.292 3.105 4.261 9.658
Resemble-Enhance 1.323 4.100 3.175 3.051 4.271 9.334
UniverSE++ 1.351 3.170 2.193 2.706 3.540 8.999
VoiceBridge 1.873 4.177 3.732 3.058 4.338 9.968
G.1 Evaluation datasets